OpenStudy (anonymous):

find the integral of -2/((t^2)+1) [0,1]

OpenStudy (anonymous):

integral of -2/((t^2)+1) is -2(arctan(t)) now u can take the limits

OpenStudy (anonymous):

thus after taking limits it is -2(arctan(1)) + 02(arctan(t)) =-2*(pi)/4 =-(pi)/2 [as arctan(1)=(pi)/4 and arctan(0)=0 ]